OpenStudy (anonymous):

Solve for x. x2-4x+2 = -2x+5

OpenStudy (anonymous):

no way to solve this exclusively for x.

OpenStudy (anonymous):

would a graph help ?

OpenStudy (anonymous):

\[x^2-4x+2+2x-5=0\] \[x^2-2x-3=0\] use the quadratic formula ... :@leealex0201 why do you think there's no way of doing it for x ???

OpenStudy (anonymous):

Never mind. I got it!! :))

OpenStudy (anonymous):

use this formula...\[\LARGE x_{1/2}\frac{-b\pm\sqrt{b^2-4ac}}{2a}\]

OpenStudy (anonymous):

we call it the x game

OpenStudy (anonymous):